solve for the linear combination of vector a &b to reach vector c on Mat lab.

vector a= [4 top -6 bottom], vector b=[2 top 6 bottom] and vector c= [14 top 6 bottom]
Answer #1

Output

4 -6 2 6 X = 2.0000 3.0000 2.000€

This means 2A+3B=C

A=[4;-6] and B=[2;6]

P=[A B] and Q=C

Script with comments

A=[4;-6] B=[2;6] C=[14;6] % we could solve the above system as system of linear equation PX=Q where % P=[AB] and Q=C, solutio

Code to copy

A=[4;-6]

B=[2;6]

C=[14;6]

% we could solve the above system as system of linear equation PX=Q where

% P=[A B] and Q=C, solution will be X=inverse(P)*Q

P=[A B]

Q=C

%pinv return inverse for an invertible matrix

X=pinv(P)*Q